Please use plain language to describe the meaning of Classless Inter-Domain Routing (CIDR) and why we need it.

Easy to understand is the key.
Answer #1

Classless Inter-Domain Routing (CIDR) is an IP addressing mechanism used as an alternative to the Classful addressing.

In classful addressing, the size of every class is fixed which leads to wastage of many IP addresses if the network needs only a few IP addresses.

CIDR overcomes the disadvantage(i.e., wastage of IP's) of Classful addressing.

CIDR is represented as:

a.b.c.d/n

where a.b.c.d represents the IP address,

n represents the no. of bits used in the network.

IP addresses in CIDR are allocated in blocks. Each block can be thought of as a box containing the IP addresses required.

CIDR follows 3 rules to allocate the IP addresses:

Rule 1: It sees that all the IP addresses allocated in a block are continuous.

Rule 2: Size of the block should be in powers of 2 (i.e., each block should contain IP's in such a way that the number of IP's can be represented as a power of 2.)

Rule 3: The first IP address of the block must be divisible by the size of the block.

Example:

Given block of IP addresses are in the range 192.168.96.32 to 192.168.96.47

Rule 1 says that the IP addresses must be continuous.

We can say that it is true as the IP's are from 192.168.96.32 to 192.168.96.47

Rule 1 is satisfied.

Rule 2 says that it should be represented in powers of 2

Number of IP addresses = 47 - 32 + 1 = 15 + 1 = 16 addresses

16 can be represented as

Rule 2 is satisfied.

Rule 3 says that the first IP address must be divisible by the size of the block.

block size = number of IP addresses = 16

First IP address is 32

32 / 16 = 2

The first IP is divisible by the size of the block.

Hence rule 3 is satisfied.

As all the rules are satisfied, the given IP addresses make up a CIDR block.

CIDR representation:

block size = number of IP addresses = 16 =

Number of bits in network part = 32 - 4 = 28

192.168.96.32/28 is the CIDR representation of the block.